i am having trouble with trying to figure out how to graph y=x-1,y=-x+3 to get (2,1)
Answer by mananth(16949)

-1 x + 1 y = -1 .............1
x=0,y=-1 (0,-1)
y=0,x=1 (1,0)
1 x + 1 y = 3 .............2
x=0, y=3 (0,3)
y=0,x=3 (3,0)
Plot these points and draw the lines. The point of intersection is the solution.

The solution is (2,1)
CHECK by solving the equation

multiply (1)by -1
Multiply (2) by 1
1 x + -1 y = 1
1 x + 1 y = 3

Add the two equations

2 x = 4

/ 2

x = 2

plug value of x in (1)

- x + y = -1
-2 + y = -1
y = -1 + 2
y = 1
y = 1